### v1

# Relativistic unitary coupled cluster method for ground-state molecular properties

Abstract
We propose a relativistic unitary coupled cluster (UCC) expectation value approach for computing first-order properties of heavy-element systems. Both perturbative (UCC3) and non-perturbative (qUCC) commutator-based formulations are applied to evaluate ground-state properties, including the permanent dipole moment (PDM), magnetic hyperfine structure (HFS) constant, and electric field gradient (EFG). The results are compared with available experimental data and those from conventional coupled cluster (CC) calculations. The non-perturbative commutator-based approach truncated at the singles and doubles level (qUCCSD) exhibits markedly better agreement with both CCSD and experiment than the perturbative UCC3 method, likely due to its improved treatment of relaxation effects.
I Introduction
A comprehensive understanding of chemical processes requires consideration of not only molecular energetics but also intrinsic properties such as dipole moments, hyperfine coupling constants, and other measurable parameters. Therefore, quantum chemical calculations of atomic and molecular properties have become a fundamental aspect of modern-day computational chemistry.[1, 2, 3, 4] However, accurate calculation of properties of atoms and molecules containing heavy elements is highly challenging due to the necessity of incorporating both relativistic and electron correlation effects in a balanced manner, especially for systems where relativistic effects are significant.[5] One of the most effective ways to account for relativistic effects in a quantum mechanical calculation for many-electronic systems is through the Dirac–Hartree–Fock (DHF) method.[6, 7] The DHF method cannot take care of the dynamic interaction of opposite-spin electrons, which is known as the electron correlation effect. Among the various electron correlation methods available, the single-reference coupled cluster (CC) method [8, 9, 10, 11, 12, 13] has emerged as one of the most accurate and systematically improvable ones. The property calculation within the CC method can be performed using two alternative approaches: one is using an expectation value approach,[14, 15] and the second is using an analytic derivative technique.[16, 17, 18, 19] It should be noted that the two approaches yield different results even for first-order properties, except in the full CC limit.[14] In the domain of relativistic coupled cluster, the calculation of first-order property has been reported[20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] using both analytic derivative and expectation value approach. The analytic derivative approach is generally more advantageous for geometrical derivatives and higher-order properties.[19] The calculation of second and higher-order properties with the framework of the relativistic CC method has been recently reported.[31, 32, 33, 34] In addition to the standard formulation of the coupled cluster method, there exist alternate ansatzes within the CC framework.[35] One can use a unitary ansatz[36] to arrive at a hermitian formulation of coupled cluster.[37, 38, 39] In recent times, the unitary coupled cluster(UCC) method has gained significant attention, not only for its computational advantages but also because of its potential applications in quantum computing.[40, 41, 42, 42] The UCC method has a particular advantage: the first-order properties calculated using both expectation value and analytic derivative will yield the same answer for a particular truncation of the cluster operator.[38] Moreover, because the standard CC energy functional is non-Hermitian, calculating first-order ground-state properties with the analytic derivative technique requires not only the CC amplitudes but also an additional set of left-vector amplitudes to ensure that the energy functional is stationary. In the UCC expectation value formalism, one can obtain the first-order properties with a single set of amplitudes. However, unlike the conventional CC theory, the UCC framework does not allow for a natural truncation of the similarity-transformed Hamiltonian. As a result, an imposed truncation is necessary. Any arbitrary truncation of the UCC functional may compromise the size-extensivity of the energy.[38] To address this challenge, Bartlett and co-workers[43] proposed a truncation scheme based on perturbative analysis of the UCC energy functional, preserving the size-extensivity of the energy. Building on this, Taube and Bartlett[44] introduced an improved truncation scheme that yields exact results for two-electron systems. However, for larger and more complex molecules, the perturbative approximation to the UCC theory(UCC(n))often fails due to the poor convergence behavior of the underlying low-order Møller-Plesset(MP) series.[45] To overcome this limitation, recent efforts have focused on truncating the UCC expansion based on the rank of the nested commutators[46, 47, 48] using Bernoulli expansion, which has shown improved performance in systems with irregular MP convergence. Furthermore, Liu et al.[48] proposed the UCCn method, which applies a perturbative truncation of the Bernoulli expansion and offers enhanced accuracy in such challenging cases. The UCC3 method has been particularly well studied[48, 49, 50, 51] in recent times for ground and excited states of small molecules. Recent work by DePrince and co-workers[45] demonstrated that the commutator-based truncation scheme achieves faster convergence toward the Full CI limit and provides more reliable results than the perturbative truncation-based approximations, especially for molecular systems away from their equilibrium geometries. In spite of many attractive features, the report of the property calculation using UCC is rather limited. Bartlett and co-workers have reported an analytic gradient technique for the UCC(4) method.[52] Sur et al. have used the relativistic UCC3 method to calculate nuclear quadrupole moment, hyperfine constants, and transition properties of atoms.[46] Dreuw and co-workers[50] have recently reported ground and excited state properties using the UCC3 method. One of the main issues with property calculations using low-order Møller–Plesset (MP) theory UCC methods is that they yield less accurate results than the standard CC approach.[52] The non-perturbative quadratic unitary coupled cluster method (qUCC) of Cheng and co-workers has been shown to give comparable performance to that of the standard CC method for energy calculations.[47] We have recently reported a relativistic variant of the qUCC method.[53] The aim of this paper is to extend the relativistic UCC method for the first-order property calculation using the expectation value approach. The structure of this paper is as follows: Section II presents the theoretical framework of the method, Section III provides the computational details, and Section IV presents the results of molecular property calculations for selected systems. Finally, conclusions are summarized in Section V.
II Theoretical framework
II.1 Relativistic Unitary Coupled Cluster Theory
The DHF method is generally the starting point for all relativistic single-reference electron correlation calculations. It extends the Hartree-Fock method by incorporating the principle of special relativity via the Dirac-Coulomb (DC) Hamiltonian,[7] , and for a molecular system under the Born-Oppenheimer approximation, can be defined as,
where denotes the momentum operator, is the rest mass, and is the speed of light. The operator represents the potential energy of the -th electron in the field of the nucleus A with a Gaussian charge distribution. The symbols and are the Dirac matrices, and denotes the identity matrix. The DHF method can be represented in matrix form
In this representation, and correspond to the large and small components of the four-component wavefunction, each of the components taking the form of a two-spinor. The operator represents the nuclear-electron interaction. Whereas and represent the Coulomb and exchange operators, respectively. The Pauli spin matrices are denoted by .
The Eq. (1) can be rewritten in the occupation number representation as
Where is the Fock matrix, defined as
and is the so-called fluctuation potential,
Here, the indices p, q, r, s represent positive-energy four-component spinors within the framework of the no-pair approximation [54]. To account for the effects of electron correlation, we employ the UCC ansatz on top of the DHF reference . In the UCC method, the ground-state wave function is expressed as
The is the excitation operator, while its Hermitian conjugate accounts for the corresponding de-excitation cluster operator. The difference forms an anti-Hermitian operator, i.e, , ensuring that the exponential operator is unitary and thus preserves the norm of the wave function. The cluster operator in the singles and doubles truncation is given by
where , , , , and , , , symbols are used to denote the occupied and virtual spinors, respectively. The ground state energy is obtained as
The cluster amplitudes in the projection based unitary coupled cluster are determined by projecting the electronic Schrödinger equation on to the excited state determinants as,
is the similarity-transformed Hamiltonian defined as,
This similarity transformed Hamiltonian can be expanded using the Baker-Campbell-Hausdorff (BCH) expansion formula as
One major challenge in UCC theory is that, unlike traditional CC theory, there is no natural finite truncation of the Baker–Campbell–Hausdorff (BCH) expansion for . This arises due to the presence of both excitation and de-excitation operator cluster operators in . Therefore, the expansion requires an imposed truncation. One needs to take special precautions to ensure the size extensivity of the truncated unitary coupled cluster method. This can be done either by applying a perturbative truncation[43] based on Møller-Plesset (MP) perturbation theory or by using a commutator-based approach[48]. The commutator-based truncation schemes using the Bernoulli numbers [55, 56] have been found to be particularly effective. It is generally derived using the superoperator (denoted by a double hat) approach[57]. A superoperator (), when applied on any arbitrary operator (), gives rise to a commutator
Using the above definition, the Eq. (12) can be written as
The above expression is equivalent to the BCH expansion Eq. (13).
Now, one can separate the using the superoperator
Where denotes the exponential Taylor series expressed as
The inverse of the above function can be expressed as
The denotes the Bernouli numbers
By left multiplying Eq. (17) with , one can arrive at
One can obtain expressions for the iterative generation of as
and
As it can be seen that the iterative generation of only includes a singly nested commutator in , which makes it advantageous over other truncation schemes to UCC. To use the iterative rule to generate the , the can be split into two parts
Where “N” denotes the non-diagonal part of the operator, which involves all pure excitation and de-excitation operators, and “R” represents the rest of the part obtained after excluding the non-diagonal part. With this definition, the UCC amplitude equations can be rewritten as
Assuming the Fock operator to be block diagonal. The iterative formula for the similarity-transformed fluctuation potential can now be written as
The above expression allows the construction of the only from the knowledge of the rest of the part of the similarity-transformed fluctuation potential. One can start with the guess and put into Eq. (26) to obtain
Which can then be plugged into Eq. (26) to obtain and so on. Therefore, the total can be expressed as
where
and so on. One can derive non-perturbative approximations to the UCC method by including different orders of in the energy and amplitude equations. Among the various approximations that the above approach offers, the qUCC method offers the best compromise between computational cost and accuracy[45]. The qUCC energy and amplitudes are as follows
The qUCC method is generally used in singles and doubles approximation (qUCCSD)[47]. Extension of the qUCCSD method to non-iterative triples correction has also been achieved[58]. Alternatively, one can derive a perturbative approximation to UCC by truncating the amplitudes in Eq. (25) in perturbation order, and it is generally denoted UCCn, where n denotes the order of perturbation. It is important to note that the UCCn approach differs from Bartlett’s UCC(n) framework[43], where the ground state energy is accurate through the order n in MP perturbation theory. In the present formulation, the amplitude equations are consistent through the n-th order.
II.2 First order Property Calculation using UCC Expectation Value Formalism
We have formulated the property calculation within the UCC method using an ADC-like intermediate state representation[59, 50]. The wave function for the Ith excited state can be defined as
Where is a CI-like excitation operator for the ground state, is a unity operator, and the ground state expectation value corresponding to an arbitrary operator () can be written as
Consequently, one can define the ground-state first-order property at the qUCCSD level
Subsequently, the excited state first-order property can be represented as
Where will result in the first-order property corresponding to excited state, and corresponds to the transition property due to excitation from to excited state. The programmable expressions for the qUCCSD first-order ground state property calculation are provided in the supporting information. One can derive the corresponding UCC3 expression by neglecting terms that contain amplitude products beyond fourth order in perturbation from Eq. (38). To assess the performance of the new UCC expectation value approach, we have calculated the permanent dipole moment (PDM), magnetic hyperfine structure (HFS) constant, and electric field gradient (EFG), and compared them with the CCSD Z-vector method and experimental results wherever available.
II.2.1 Permanent Dipole Moment (PDM)
The PDM operator () of a molecular system can be expressed as the sum of the electronic contribution (EC) and the nuclear contribution (NC),
In the above equation, e is the charge of the electron, is the position vector of the electron from the origin, is the atomic number and is the position vector of nucleous. Now eq (37) can be re-written as
The electronic contribution to in Eq. (41) depends on the choice of correlation method and basis set, whereas the nuclear part is independent of both the correlation method and the basis set; it remains constant for a particular molecule with a certain charge and at a specific bond length.
II.2.2 Hyperfine Structure (HFS) Constant
The magnetic hyperfine splitting in atoms arises from the interaction between the electromagnetic field generated by the electrons and the magnetic dipole moment of the nucleus. Within the magnetic dipole approximation, the vector potential at the position of electron due to the nuclear magnetic moment is given by
where is the position vector of the electron relative to the nucleus.In a relativistic framework, the corresponding magnetic hyperfine interaction for all electron can be written as
where denotes the Dirac matrices of electron . For a diatomic molecule, the magnetic hyperfine interaction can be separated into parallel and perpendicular components of the magnetic HFS constant. The z projection of the expectation value of the corresponding HFS operator gives the while the x/y projections give the constant, expressed as
Where denotes the nuclear spin quantum number and is the projection of total electronic angular momentum along the z axes of the diatomic molecule.
II.2.3 Electric Field Gradient (EFG)
The EFG is a traceless symmetric second-rank tensor that describes the spatial variation of the electric field around a nucleus due to the distribution of surrounding charges. The EFG provides valuable information about the local electronic structure and symmetry of atoms or molecules, particularly in cases where the environment of the nucleus is asymmetric. Within the principal axes coordinate system, the expectation value of the component of EFG at the position () of nucleus K is defined as
The expectation value of the above operator will lead to two terms
The first term in the above equation is the electronic contribution, named , which depends on the ground state wave-function. The second term , which arises from the other nuclei in the molecule, depends only on the position of that remaining nucleus and is independent of the electronic wave function.
III Computational and implementation details
The four-component UCC3 and qUCCSD expectation value approach for the calculation of the first-order property has been implemented in our in-house quantum chemistry software package, BAGH.[60]
It is primarily developed in Python, with computationally intensive parts written in Fortran and Cython. BAGH relies on external software packages to generate one and two-electron integrals, and it is interfaced with PySCF[61, 62, 63], socutils[64], DIRAC[65], and GAMESS-US[66]. The DIRAC package is used to solve the DHF equations and to compute one- and two-electron integrals, along with one-electron property integrals for the PDM and HFS calculations. For the EFG calculation, the DHF orbital and the necessary integrals are instead generated using the PySCF package[61, 62, 63] for this work.
Dipole moment calculations were performed for a series of alkaline earth metal monofluoride (AF, where A = Mg, Ca, Sr, Ba), employing a quadruple-zeta basis set, Dyall.cv4z[67, 68] for the metal atoms and cc-pCVQZ[69] for the fluorine atom.
For HFS constant calculations, the aug-cc-pCV5Z[69] basis set was used for the hydrogen atom, aug-cc-pCVQZ for F, Be, and Mg atoms, and Dyall.cv4z for the Ca atom. For the MgH and MgF, a virtual orbital energy cutoff of 10 a.u. was applied, while a cutoff of 15 a.u. was used for CaH and MgF.
EFG calculations were carried out using Dyall.cv2z[70] basis set for all FX [where, X= F, Cl, Br, I, At] molecules.
IV Result and Discussion
IV.1 Permanent Dipole Moment
Table 1 summarizes the PDM of the Group II monofluoride molecules (MgF, CaF, SrF, and BaF) in UCC3, qUCCSD, and CCSD Z-vector methods, along with their available experimental data. All the calculated values are presented in Debye units. To ensure consistency with the Z-vector method, we computed our results using the same molecular geometry, basis set, and cut-off energy as those employed by Pal and co-workers.[71]
Figure 1(a) presents the comparison of UCC3, the qUCCSD method, with the CCSD Z-vector method, and it can be seen that the qUCCSD results show an excellent agreement with the CCSD Z-vector approach. Whereas the deviation in the UCC3 method is quite high. This discrepancy arises because the qUCCSD method contains a more complete treatment of the single excitation (up to quadratic term in the amplitudes equation), whereas they are linear in the UCC3 amplitude equation. For the MgF molecule, both UCC3 and qUCCSD approaches agree remarkably with each other, but the agreement deteriorates as we go down the group. In the case of CaF, the UCC3 method shows a deviation of 0.22 Debye from the CCSD Z-vector results, while the deviation from the CCSD Z-vector results is 0.01 Debye for the qUCCSD method. Similarly, for SrF, the deviations from the CCSD Z-vector method are 0.39 Debye and 0.03 Debye for UCC3 and qUCCSD, respectively. The BaF molecule shows the largest deviation for UCC3 at 0.82 Debye, whereas the qUCCSD method maintains high accuracy with a deviation of just 0.01 Debye. These findings demonstrate that the qUCCSD method yields results in excellent agreement with those obtained using the Z-vector approach.
We have compared our results with experimental measurements for CaF, SrF, and BaF. No experimental measurements are available for MgF. To assess the accuracy of our results, we compared them with available experimental data, with deviations expressed as percentage errors, denoted by .
Figure 1(b) presents a comparison of from the experimental results for UCC3, qUCCSD, and CCSD Z-vector methods. It can be seen that the UCC3 method displays a larger deviation from the experimental values, whereas the qUCCSD and Z-vector methods yield more accurate results and are generally very close to each other. The UCC3 method shows the smallest deviation with respect to the experiment for CaF and is 8.15 . The qUCCSD and Z-vector methods yield much lower deviations of 1.70 and 1.30, respectively, and it is observed for SrF. The UCC3 shows the highest deviation of 22.63 for BaF. The qUCCSD and Z-vector results show very good agreement with the experiment for BaF with remarkably improved values of 3.36, and 3.11, respectively.
| Molecule | UCC3 | qUCCSD | Z-vector[71] | Expt. | |||
| UCC3 | qUCCSD | Z-vector | |||||
| MgF | -3.13 | -3.14 | -3.10 | – | – | – | – |
| CaF | -3.33 | -3.12 | -3.11 | 3.07[72] | 8.31 | 1.70 | 1.30 |
| SrF | -3.75 | -3.39 | -3.36 | 3.4676[73] | 8.15 | 2.54 | 3.10 |
| BaF | -3.90 | -3.07 | -3.08 | 3.179[74] | 22.63 | 3.36 | 3.11 |
IV.2 Hyperfine Structure Constant
The HFS constants are sensitive to the quality of the wave function in the nuclear region and can be a good test for the quality of the wave function obtained from a particular method. For this purpose, we have computed the HFS constants for a series of diatomic molecules, including BeH, MgH, CaH, and CaF, using both the UCC3 and qUCCSD methods. The computed parallel () and perpendicular () components of the HFS constants are presented in Tables 2 and 3 , respectively. To ensure consistency, our results are compared with those from the Z-vector method, using identical basis sets, geometries, and cutoff values as described in the reference.[75] The data presented in Tables 2 and 3 reveal that the qUCCSD results show better agreement with both Z-vector and the experimental values compared to the UCC3 results, which exhibit more deviation with respect to the experiment. The maximum and minimum deviation of UCC3 and qUCCSD for values from the Z-vector results are observed in different systems. For UCC3, the largest deviation (92.87 MHz) from the CCSD Z vector result occurs in 1H in MgH, while the smallest (0.82 MHz) is found for 25Mg in MgF. In contrast, qUCCSD shows a maximum deviation of 6.61 MHz for 19F in MgF and a minimum of 0.89 MHz for 1H in CaH.
In the case of , the maximum deviation of 93.64 MHz from the CCSD Z-vector method is obtained in UCC3 again for 1H in MgH. The minimum deviation for UCC3 (0.63 MHz) is observed for 25Mg in MgF. The maximum deviation in the qUCCSD method from the CCSD Z-vector is 4.98 MHz observed for 25Mg in MgF. Figures 3 (a) and 3 (b) present a comparison of the and values, respectively, obtained using UCC3, qUCCSD and Z-vector method. From the above observation, the results clearly demonstrate that the qUCCSD approach consistently yields better agreement with the CCSD Z-vector results compared to the UCC3 method. Our calculated results have also been compared with the available experimental data. In figures 3(a) and 3(b) represent the percentage deviation () of the and values, respectively, obtained using UCC3, qUCCSD, and Z-vector method. For , the largest in UCC3 occurs at 1H in CaH, which 51.34 relative deviation from experimental. The higher relative deviation observed in the qUCCSD and CCSD Z-vector methods is 17.42 MHz and 16.77 MHz, respectively, and it is observed for the 25 Mg nucleus in MgH. In the case of the component, the highest deviation is observed for the 1H nucleus in the CaH molecule, and it shows a 51.22 with the UCC3 method. The qUCCSD approach shows its highest deviation of 23.78 for the 19F nucleus in CaF, which is close to the corresponding CCSD Z-vector value of 20.94 . From the overall trends observed in the table, it reveals that the qUCCSD method demonstrates very good agreement with experimental values for both and of the hyperfine coupling constant, showing close agreement with the corresponding Z-vector results. The UCC3 method shows significant deviation with respect to experiments in some cases.
| Molecule | Atom | UCC3 | qUCCSD | Z-vector[75] | Expt. | |||
| UCC3 | qUCCSD | Z-vector | ||||||
| BeH | 1H | 227.02 | 201.94 | 199.5 | 201[76] | 12.94 | 0.47 | 0.75 |
| 9Be | -195.21 | -200.35 | -201.3 | -208[76] | 6.15 | 3.68 | 3.22 | |
| MgH | 1H | 369.07 | 281.12 | 276.2 | 298[77] | 23.85 | 5.66 | 7.31 |
| 25Mg | -169.97 | -186.64 | -188.1 | -226[77] | 24.79 | 17.42 | 16.77 | |
| CaH | 1H | 208.84 | 142.81 | 143.7 | 138[77] | 51.34 | 3.49 | 4.31 |
| 43Ca | -321.28 | -327.40 | -331.2 | – | – | – | – | |
| MgF | 19F | 347.41 | 315.99 | 322.6 | 331[78] | 4.96 | 4.53 | 2.54 |
| 25Mg | -280.98 | -283.02 | -281.8 | – | – | – | – | |
| CaF | 19F | 127.05 | 129.11 | 131.7 | 149[78] | 14.73 | 13.35 | 11.61 |
| 43Ca | -433.92 | -421.60 | -426.4 | – | – | – | – | |
| Molecule | Atom | UCC3 | qUCCSD | Z-vector[75] | Expt. | |||
| UCC3 | qUCCSD | Z-vector | ||||||
| BeH | 1H | 208.90 | 183.43 | 181.0 | 190.8[76] | 9.49 | 3.86 | 5.14 |
| 9Be | -180.61 | -185.83 | -186.8 | -194.8[76] | 7.28 | 4.60 | 4.11 | |
| MgH | 1H | 364.74 | 276.08 | 271.1 | 264[77] | 38.16 | 4.58 | 2.69 |
| 25Mg | -154.15 | -170.61 | -172.0 | -218.0[77] | 29.29 | 21.74 | 21.10 | |
| CaH | 1H | 202.64 | 138.22 | 139 | 134[77] | 51.22 | 3.15 | 3.73 |
| 43Ca | -293.97 | -301.46 | -305.2 | – | – | – | – | |
| MgF | 19F | 128.23 | 148.59 | 148.7 | 143[78] | 10.33 | 3.91 | 3.99 |
| 25Mg | -268.97 | -270.82 | -269.6 | – | – | – | – | |
| CaF | 19F | 62.51 | 80.79 | 83.8 | 106[78] | 41.03 | 23.78 | 20.94 |
| 43Ca | -415.89 | -403.98 | -408.9 | – | – | – | – | |
IV.3 Electric Field Gradient
To further assess the performance of the newly developed UCC methods, we have calculated the electronic component of the EFG value () for the series of dihalogen molecules FY (Y = F, Cl, Br, I, At) and compared them with the four-component CCSD Z-vector results. The reference CCSD Z-vector values were adopted from the work of Aucar et al.[79] To maintain consistency, all calculations in this work were carried out using the same dyall.cv2z basis set and molecular geometries as employed in the ref [79]. The molecular geometries for the series of dihalogen molecules are also provided in the supporting information. Figure 4(a) provides a comparison of the UCC3 and qUCCSD values with the reference CCSD Z-vector method for the zz-component (with the z axis chosen along the molecular bond) of the electronic component of EFG values () for atoms in the FY series. The corresponding values are provided in Table 4. It can be seen that the qUCCSD values closely follow the CCSD values throughout the data set, with the bars almost overlapping in most cases. The UCC3 method displays small but noticeable deviations, particularly for heavier atoms such as I and At. Table 4 also presents their deviations from the reference values, which in this case are CCSD Z-vector values, as no experimental values are available. From the table, it is clear that qUCCSD values show smaller deviations from CCSD results compared to UCC3, often by an order of magnitude. The highest deviation in UCC3 is observed to be 1.008 a.u., which is observed for At in FAt. The qUCCSD method, on the other hand, shows a deviation of 0.086 a.u. for the electronic component of EFG values ().
| Molecule | Atom | UCC3 | qUCCSD | Z-vector[79] | ||
| UCC3 | qUCCSD | |||||
| F2 | F | 5.145 | 5.174 | 5.164 | 0.364 | 0.194 |
| FCl | F | 3.013 | 2.974 | 2.967 | 1.554 | 0.220 |
| Cl | 6.850 | 6.918 | 6.933 | 1.196 | 0.221 | |
| FBr | F | 1.872 | 1.773 | 1.768 | 5.892 | 0.280 |
| Br | 14.363 | 14.607 | 14.622 | 1.769 | 0.105 | |
| IF | F | 0.810 | 0.677 | 0.683 | 18.563 | 0.933 |
| I | 20.233 | 20.85 | 20.857 | 2.991 | 0.015 | |
| FAt | F | -0.251 | -0.531 | -0.501 | 49.936 | 5.988 |
| At | 36.803 | 37.897 | 37.811 | 2.666 | 0.227 | |
V Conclusion
We present the theory, implementation, and benchmarking of the first-order property calculation using the unitary coupled cluster expectation value method. The Hermitian formulation within the qUCCSD method allows a simple definition of the first-order property using the expectation value approach and does not require the calculation of an additional set of amplitudes, as in the standard coupled cluster Z-vector method. We have calculated the permanent dipole moment, hyperfine structure constant, and electric field gradient of a series of molecules containing heavy atoms. The calculated values using the qUCCSD expectation value approach are in close agreement with the CCSD Z-vector results and are consistent with the available experimental values. However, the perturbative UCC3 method does not give consistent results and often shows large errors. The cause is presumably the incomplete treatment of the single excitation in the UCC3 level. The present work shows that the UCC method gives an attractive option for the calculation of first-order properties even on a classical computer. It would be interesting to extend the unitary coupled cluster method to second and higher-order properties. Work is in progress in that direction.
Supplementary Material
The Supplementary Material contains the programmable expressions for the ground state first-order property in the UCC3 and qUCCSD methods and the molecular geometry for the series of dihalogen molecules.
Acknowledgements.
The authors acknowledge the support from the IIT Bombay, DST-SERB CRG (Project No. CRG/2022/005672) and MATRICS (Project No. MTR/2021/000420) projects, and IIT Bombay supercomputational facility and C-DAC supercomputing resources (PARAM Smriti and PARAM Brahma) for computational time. KM acknowledges CSIR-HRDG for a senior research fellowship. SC acknowledges the Prime Minister’s Research Fellowship (PMRF). The authors acknowledge Juan J. Aucar (Institute for Modelling and Innovative Technology, IMIT (CONICET-UNNE), Argentina) for his valuable feedback and discussions during the preparation of this manuscript.References
- Helgaker et al. [2012] T. Helgaker, S. Coriani, P. Jørgensen, K. Kristensen, J. Olsen, and K. Ruud, Chem. Rev. 112, 543 (2012).
- Neese [2009] F. Neese, Coord. Chem. Rev. 253, 526 (2009).
- Wilson and Diercksen [2013] S. Wilson and G. H. Diercksen, Methods in computational molecular physics, Vol. 293 (Springer Science & Business Media, 2013).
- Norman, Ruud, and Saue [2018] P. Norman, K. Ruud, and T. Saue, Principles and practices of molecular properties: Theory, modeling, and simulations (John Wiley & Sons, 2018).
- Visscher, Lee, and Dyall [1996a] L. Visscher, T. J. Lee, and K. G. Dyall, J. Chem. Phys. 105, 8769 (1996a).
- Dyall and Fægri Jr. [2007] K. G. Dyall and K. Fægri Jr., Introduction to Relativistic Quantum Chemistry (Oxford University Press, Oxford, New York, 2007).
- Reiher and Wolf [2014] M. Reiher and A. Wolf, Relativistic quantum chemistry: the fundamental theory of molecular science (John Wiley & Sons, 2014).
- Cizek [1967] J. Cizek, “Advances in chemical physics: correlation effects in atoms and molecules,” (1967).
- Mukherjee and Pal [1989] D. Mukherjee and S. Pal, Adv. Quantum Chem. 20, 291 (1989).
- Bartlett and Purvis [1978] R. J. Bartlett and G. D. Purvis, Int. J. Quantum Chem. 14, 561 (1978).
- Bartlett and Musiał [2007] R. J. Bartlett and M. Musiał, Rev. Mod. Phys. 79, 291 (2007).
- Crawford and Schaefer III [2007] T. D. Crawford and H. F. Schaefer III, Rev. Comput. Chem. 14, 33 (2007).
- Shavitt and Bartlett [2009] I. Shavitt and R. J. Bartlett, Many-body methods in chemistry and physics: MBPT and coupled-cluster theory (Cambridge university press, 2009).
- Pal [1984a] S. Pal, Theor. Chim. Acta 66, 151 (1984a).
- Bartlett and Noga [1988] R. J. Bartlett and J. Noga, Chem. Phys. Lett. 150, 29 (1988).
- Monkhorst [1977] H. J. Monkhorst, Int. J. Quantum Chem. 12, 421 (1977).
- Salter, Trucks, and Bartlett [1989] E. A. Salter, G. W. Trucks, and R. J. Bartlett, J. Chem. Phys. 90, 1752 (1989).
- Scheiner et al. [1987] A. C. Scheiner, G. E. Scuseria, J. E. Rice, T. J. Lee, and I. Schaefer, Henry F., J. Chem. Phys. 87, 5361 (1987).
- Koch and Jo/rgensen [1990] H. Koch and P. Jo/rgensen, J. Chem. Phys. 93, 3333 (1990).
- Visscher, Dyall, and Lee [1995] L. Visscher, K. G. Dyall, and T. J. Lee, Int. J. Quantum Chem. 56, 411 (1995).
- Visscher, Lee, and Dyall [1996b] L. Visscher, T. J. Lee, and K. G. Dyall, J. Chem. Phys. 105, 8769 (1996b).
- Petrov et al. [2002] A. Petrov, N. Mosyagin, T. Isaev, A. Titov, V. Ezhov, E. Eliav, and U. Kaldor, Phys. Rev. Lett. 88, 073001 (2002).
- Eliav, Kaldor, and Ishikawa [1994a] E. Eliav, U. Kaldor, and Y. Ishikawa, Phys. Rev. A 49, 1724 (1994a).
- Eliav, Kaldor, and Ishikawa [1994b] E. Eliav, U. Kaldor, and Y. Ishikawa, Phys. Rev. A 50, 1121 (1994b).
- Eliav, Kaldor, and Ishikawa [1995] E. Eliav, U. Kaldor, and Y. Ishikawa, Phys. Rev. A 51, 225 (1995).
- Shee, Visscher, and Saue [2016] A. Shee, L. Visscher, and T. Saue, J. Chem. Phys. 145 (2016).
- Sasmal et al. [2015a] S. Sasmal, H. Pathak, M. K. Nayak, N. Vaval, and S. Pal, Phys. Rev. A 91, 022512 (2015a).
- Prasannaa et al. [2016] V. Prasannaa, S. Sreerekha, M. Abe, V. Bannur, and B. Das, Phys. Rev. A 93, 042504 (2016).
- Sasmal et al. [2015b] S. Sasmal, H. Pathak, M. K. Nayak, N. Vaval, and S. Pal, Phys. Rev. A 91, 030503 (2015b).
- Das and Mukherjee [2005] B. Das and D. Mukherjee, Eur. Phys. J. D 32, 25 (2005).
- Yuan et al. [2024] X. Yuan, L. Halbert, J. V. Pototschnig, A. Papadopoulos, S. Coriani, L. Visscher, and A. S. Pereira Gomes, J. Chem. Theory Comput. 20, 677 (2024).
- Yuan et al. [2023] X. Yuan, L. Halbert, L. Visscher, and A. S. Pereira Gomes, J. Chem. Theory Comput. 19, 9248 (2023).
- Chakraborty, Mukhopadhyay, and Dutta [2025] S. Chakraborty, T. Mukhopadhyay, and A. K. Dutta, J. Phys. Chem. A (2025), 10.1021/acs.jpca.4c03584, pMID: 40152233.
- Chakraborty et al. [2025] S. Chakraborty, A. Manna, T. D. Crawford, and A. K. Dutta, J. Chem. Phys. 163, 044111 (2025).
- Szalay, Nooijen, and Bartlett [1995] P. G. Szalay, M. Nooijen, and R. J. Bartlett, J. Chem. Phys. 103, 281 (1995).
- Kutzelnigg [1977] W. Kutzelnigg, in Methods of Electronic Structure Theory, edited by H. F. Schaefer (Springer US, Boston, MA, 1977) pp. 129–188.
- Bartlett, Kucharski, and Noga [1989a] R. J. Bartlett, S. A. Kucharski, and J. Noga, Chem. Phys. Lett. 155, 133 (1989a).
- Pal [1984b] S. Pal, Theor. Chem. Acc. 66, 207 (1984b).
- Hoffmann and Simons [1988] M. R. Hoffmann and J. Simons, J. Chem. Phys. 88, 993 (1988).
- McArdle et al. [2020] S. McArdle, S. Endo, A. Aspuru-Guzik, S. C. Benjamin, and X. Yuan, Rev. Mod. Phys. 92, 015003 (2020).
- Cao et al. [2019] Y. Cao, J. Romero, J. P. Olson, M. Degroote, P. D. Johnson, M. Kieferová, I. D. Kivlichan, T. Menke, B. Peropadre, N. P. Sawaya, et al., Chem. Rev. 119, 10856 (2019).
- Bauer et al. [2020] B. Bauer, S. Bravyi, M. Motta, and G. K.-L. Chan, Chem. Rev. 120, 12685 (2020).
- Bartlett, Kucharski, and Noga [1989b] R. J. Bartlett, S. A. Kucharski, and J. Noga, Chem. Phys. Lett. 155, 133 (1989b).
- Taube and Bartlett [2006] A. G. Taube and R. J. Bartlett, Int. J. Quantum Chem. 106, 3393 (2006).
- Phillips et al. [2025] J. T. Phillips, L. N. Koulias, S. H. Yuwono, and A. E. DePrince III, Mol. Phys. , e2522382 (2025).
- Sur et al. [2008] C. Sur, R. K. Chaudhuri, B. K. Sahoo, B. Das, and D. Mukherjee, J. Phys. B: At. Mol. Opt. Phys. 41, 065001 (2008).
- Liu and Cheng [2021] J. Liu and L. Cheng, J. Chem. Phys. 155 (2021).
- Liu et al. [2018] J. Liu, A. Asthana, L. Cheng, and D. Mukherjee, J. Chem. Phys. 148 (2018).
- Hodecker et al. [2020] M. Hodecker, S. M. Thielen, J. Liu, D. R. Rehn, and A. Dreuw, J. Chem. Theory Comput. 16, 3654 (2020), pMID: 32396348, https://doi.org/10.1021/acs.jctc.0c00335 .
- Hodecker and Dreuw [2020] M. Hodecker and A. Dreuw, J. Chem. Phys. 153, 084112 (2020).
- Dempwolff, Hodecker, and Dreuw [2022] A. L. Dempwolff, M. Hodecker, and A. Dreuw, J. Chem. Phys. 156, 054114 (2022).
- Watts, Trucks, and Bartlett [1989] J. D. Watts, G. W. Trucks, and R. J. Bartlett, Chem. Phys. Lett. 157, 359 (1989).
- Majee et al. [2024] K. Majee, S. Chakraborty, T. Mukhopadhyay, M. K. Nayak, and A. K. Dutta, J. Chem. Phys. 161 (2024).
- Sucher [1980] J. Sucher, Phys. Rev. A 22, 348 (1980).
- Kutzelnigg and Koch [1983] W. Kutzelnigg and S. Koch, J. Chem. Phys. 79, 4315 (1983).
- Kutzelnigg [1982] W. Kutzelnigg, J.Chem. Phys. 77, 3081 (1982).
- Prasad, Pal, and Mukherjee [1980] M. D. Prasad, S. Pal, and D. Mukherjee, Pramana 15, 531 (1980).
- Majee et al. [2025] K. Majee, J. Šimunek, J. Noga, and A. K. Dutta, “A perturbative triples correction to relativistic quadratic unitary coupled cluster method: Theory, implementation and benchmarking,” (2025), arXiv:2509.03634 [physics.chem-ph] .
- Schirmer [1991] J. Schirmer, Phys. Rev. A 43, 4647 (1991).
- Dutta et al. [2023] A. Dutta, A. Manna, B. Jangid, K. Majee, K. Surjuse, M. Mukherjee, M. Thapa, S. Arora, S. Chamoli, S. Haldar, et al., “Bagh: A quantum chemistry software package,” (2023).
- Sun et al. [2020] Q. Sun, X. Zhang, S. Banerjee, P. Bao, M. Barbry, N. S. Blunt, N. A. Bogdanov, G. H. Booth, J. Chen, Z.-H. Cui, et al., J. Chem. Phys. 153 (2020).
- Sun et al. [2018] Q. Sun, T. C. Berkelbach, N. S. Blunt, G. H. Booth, S. Guo, Z. Li, J. Liu, J. D. McClain, E. R. Sayfutyarova, S. Sharma, et al., Wiley Interdiscip. Rev. Comput. Mol. Sci. 8, e1340 (2018).
- Sun [2015] Q. Sun, J. Comput. Chem. 36, 1664 (2015).
- Xubwa [2024] Xubwa, “socutils,” (2024), accessed: 2024-12-24.
- Jensen et al. [2022] H. A. Jensen, R. Bast, A. Gomes, T. Saue, L. Visscher, I. Aucar, V. Bakken, C. Chibueze, J. Creutzberg, and K. Dyall, (2022).
- Barca et al. [2020] G. M. Barca, C. Bertoni, L. Carrington, D. Datta, N. De Silva, J. E. Deustua, D. G. Fedorov, J. R. Gour, A. O. Gunina, E. Guidez, et al., J. Chem. Phys. 152 (2020).
- Dyall [2009] K. G. Dyall, J. Phys. Chem. A 113, 12638 (2009).
- Dyall [2016] K. G. Dyall, Theor. Chim. Acta 135, 128 (2016).
- Dunning Jr [1989] T. H. Dunning Jr, J. Chem. Phys. 90, 1007 (1989).
- Dyall [2011] K. G. Dyall, Theor. Chim. Acta 129, 603 (2011).
- Haldar et al. [2021] S. Haldar, K. Talukdar, M. K. Nayak, and S. Pal, Int. J. Quantum Chem. 121, e26764 (2021).
- Anand et al. [2022] A. Anand, P. Schleich, S. Alperin-Lea, P. W. Jensen, S. Sim, M. Díaz-Tinoco, J. S. Kottmann, M. Degroote, A. F. Izmaylov, and A. Aspuru-Guzik, Chem. Soc. Rev. 51, 1659 (2022).
- Ernst et al. [1985] W. Ernst, J. Kändler, S. Kindt, and T. Törring, Chem. Phys. Lett. 113, 351 (1985).
- Ernst, Kändler, and Törring [1986] W. Ernst, J. Kändler, and T. Törring, J. Chem. Phys. 84, 4769 (1986).
- Sasmal et al. [2016] S. Sasmal, K. Talukdar, M. K. Nayak, N. Vaval, and S. Pal, J. Chem. Sci. 128, 1671 (2016).
- Knight Jr, Brom Jr, and Weltner Jr [1972] L. Knight Jr, J. Brom Jr, and W. Weltner Jr, J. Chem. Phys. 56, 1152 (1972).
- Knight Jr and Weltner Jr [1971] L. B. Knight Jr and W. Weltner Jr, J. Chem. Phys. 54, 3875 (1971).
- Knight Jr et al. [1971] L. Knight Jr, W. Easley, W. Weltner Jr, and M. Wilson, J. Chem. Phys. 54, 322 (1971).
- Aucar, Maldonado, and Melo [2021] J. J. Aucar, A. F. Maldonado, and J. I. Melo, Int. J. Quantum Chem. 121, e26769 (2021).